OpenStudy (lgg23):

Solve for x in terms of k. log base 5 of x - log base 5 of (x+10) = log base 5 of k

OpenStudy (anonymous):

log5(x)-log5(x+10) = log5(k) log5(x/x+10) = log5(k) x/(x+10) = k x=k(x+10) x=kx+10k x(1-k)=10k x=10k/(1-k)

OpenStudy (lgg23):

how did you get the second to last step? where its x(1-k)=10k

OpenStudy (anonymous):

x=kx+10k group like terms i.e. terms with x x-kx=10k x(1-k)=10k x=10k/(1-k)

OpenStudy (lgg23):

you went through the problem faster then i could process it in my head. thank you.